OpenStudy (anonymous):

Find the x and y intercept of the circle x^2-10x+y^2+2x=-17

OpenStudy (anonymous):

I meant x^2-10x+y^2+2y=-17

OpenStudy (jdoe0001):

you know what a "perfect square trinomial" is, right?

OpenStudy (anonymous):

(x-5)^2+(y+1)^2=9 I got that far, but I ran into imaginary numbers when substituting x for 0 to find the y-intercept.

OpenStudy (jdoe0001):

well, let's just start by grouping only and then we'll "complete the square" by using our very good friend, Mr Zero \(\bf x^2-10x+y^2+2y=-17 \implies (x^2-10x)+(y^2+2y)=-17 \) what would be the 3rd number in each grouped binomial, to get us a "perfect square trinomial"?

OpenStudy (anonymous):

25 and 1, respectively. 25+1=26 Add the same number to both sides, which results in agjnpggg=9

OpenStudy (anonymous):

quantity, rather.

OpenStudy (jdoe0001):

yeap :)

OpenStudy (jdoe0001):

you add and subtract, we're just borrowing from 0, so +somevalue -somevalue = 0

OpenStudy (jdoe0001):

ohhh... scratch that

OpenStudy (jdoe0001):

\(\bf x^2-10x+y^2+2y=-17 \implies (x^2-10x)+(y^2+2y)=-17\\ (x^2-10x+5^2)+(y^2+2y+1^2)=-17 \\\implies (x-5)^2+(y+1)^2 - 5^2-1^2 = 17\\\quad \\ (x-5)^2+(y+1)^2 = 9 \implies (x-5)^2+(y+1)^2 = 3^2\)

OpenStudy (anonymous):

Yep.

OpenStudy (jdoe0001):

hmm, anyhow, I missed a -17 there, ....\(\bf x^2-10x+y^2+2y=-17 \implies (x^2-10x)+(y^2+2y)=-17\\ (x^2-10x+5^2)+(y^2+2y+1^2)=-17 \implies (x-5)^2+(y+1)^2 - 5^2-1^2 = -17\\\quad \\ (x-5)^2+(y+1)^2 = 9 \implies (x-5)^2+(y+1)^2 = 3^2 \)

OpenStudy (jdoe0001):

so that's the circle's center and radius

OpenStudy (anonymous):

(5,-1)=Center 3=Radius. Yup.

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

But x and y intercepts ;-;

OpenStudy (jdoe0001):

well, to get the "x" intercept, set y =0, solve for "x" to get the "y" intercept, set x =0, solve for "y"

OpenStudy (jdoe0001):

... I guess I didn't read careful enough... though you needed the equation, anyhow, it's easier to get the intercepts with it

OpenStudy (anonymous):

(x−5)^2+(y+1)^2=9 (0−5)^2+(y+1)^2=9 25+(y+1)^2=9 (y+1)^2=-16 Now what?

OpenStudy (jdoe0001):

hmm, well... that will give an imaginary value

OpenStudy (anonymous):

I think you would reject it. (x−5)^2+(0+1)^2=9 (x−5)^2=8 x=sqrt(8)+5 But it's really ugly, so I didn't like it.

OpenStudy (jdoe0001):

notice the center of the circle, (5, -1)  meaning, the circle never touches the y-axis, thus the imaginary value for the y-intercept

OpenStudy (jdoe0001):

-->But it's really ugly, so I didn't like it. <---- hehhe well, it doesn't have to be good-looking, it just have to be valid =)

OpenStudy (anonymous):

Right, then. Thanks.